### v1

@@ 3 Results
In this section we present our results for the finite-quark-mass effects in the gluon-gluon-fusion channel for various computational setups. Following the Higgs Working Group recommendations LHCHiggsCrossSectionWorkingGroup:2016ypw , our reference computation uses the 5FS, an OS-renormalised top quark with a pole mass of $m_{t}=173.06$ GeV [^2] The Higgs-Working-Group recommends using an OS-mass of $m_{t}=172.4$. The different value was chosen because of the rationalisation procedure described in sec. 2. We have verified, that choosing a different mass has a negligible impact on the final result. and an $\overline{\text{MS}}$-renormalised bottom mass of $\overline{m}_{b}(\overline{m}_{b})=4.18$ GeV. We then explore the impact of alternative mass-renormalisation schemes for both the top and the bottom quark. In addition, we perform the computation in the 4FS to examine if the flavour schemes are compatible.
@@
Table 3: Effect of the finite top-quark mass on the gluon-gluon fusion cross section for two different computational setups. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation.
@@
Tab. 5 presents the results for the top-bottom interference contribution to the cross section $\sigma_{t\times b}$. As already noticed in ref. Czakon:2023kqm , the interference converges quite badly if the bottom-quark mass is OS-renormalised, as indicated by the alternating perturbative corrections of a similar order of magnitude (see third column of tab. 5). Furthermore, scale uncertainties fail to correctly estimate the effects of higher orders and actually increase from NLO to NNLO, reaching 15%. However, results for $\overline{\text{MS}}$-renormalised bottom-quark masses (second column) show a much better perturbative convergence and additionally, we observe a decrease in scale uncertainties to below 8%. Adapting a mixed renormalisation scheme, in which only the Yukawa coupling is renormalised in the $\overline{\text{MS}}$ scheme, while the mass in the propagators is kept in the OS scheme, also results in a better convergence and reduced scale uncertainties as noted in ref. Czakon:2023kqm . Without the inclusion of electroweak corrections this mixed approach is valid, but if these effects are taken into account as well, the Higgs vacuum expectation value acquires an additional finite renormalisation that propagates to the masses of gauge bosons. Since the masses of the gauge bosons are defined in the on-shell scheme by default, the mixed approach is inconsistent. Ignoring this issue, we note that the main improvements to the perturbation series of our problem are already captured by the mixed scheme. For the top-quark mass, the renormalisation scheme seems to have very little impact on the final result. Lastly, the NNLO result in the 4FS is 5.2% lower than in the 5FS and agrees well within scale uncertainties. The magnitude of this decrease is of a similar size as observed in HEFT (see tab. 2).
Table 5: Top-bottom interference contribution to the gluon-gluon fusion cross section for various computational setups. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation.
@@
Figure 6: Interference contribution to the Higgs-$p_{T}$ (left) and Higgs-rapidity (right) distribution. Compared are the results for an $\overline{\text{MS}}$- and OS-renormalised bottom-quark mass. The lower panel shows the ratio to the distribution including bottom- and top-quark effects at NNLO. Coloured bands indicate the envelope of the seven-point scale variation and the error bars represent the Monte Carlo uncertainties.
@@
Figure 7: Higgs-$p_{T}$ (left) and Higgs-rapidity (right) distribution for the gluon-gluon fusion channel. Compared are results obtained in HEFT with those including finite-quark-mass effects in the square of the top-quark-loop amplitude and in the interference of the top-quark-loop and bottom-quark-loop amplitudes. The bottom-quark mass is renormalised in the $\overline{\text{MS}}$ scheme, while the top-quark mass is renormalised in the OS scheme.
For the Higgs-rapidity distribution in fig. 7 we also observe that the scale-uncertainty bands of the different pertubative orders do not overlap. For both full QCD and HEFT, the difference between NLO and NNLO amounts to a rapidity independent upwards shift of about 20%. Similarly, finite quark masses also only shift the distribution, this time by around $-4\%$, an effect which is directly derived from the interference contribution in fig. 6.
@@ 4 Conclusions
In this work, we studied the impact of using variable flavour number and different mass-renormalisation schemes on the Higgs production cross section in the gluon-gluon fusion channel. We found that the two different flavour schemes are in excellent agreement, finally justifying our treatment of the bottom-quark mass in the 5FS. Our computation in two different renormalisation schemes showed that renormalising the bottom-quark mass in the $\overline{\text{MS}}$ scheme results in significantly smaller scale uncertainties and displays a better perturbative convergence. The renormalisation of the top-quark mass on the other hand does not impact the cross section significantly. We therefore recommend using the 5FS, an $\overline{\text{MS}}$-renormalised bottom mass and an OS-renormalised top mass going forward. The cross section for this set-up reads [^4] The $\mathrm{N}^{3}\mathrm{LO}$ HEFT cross section was computed using SusHi Harlander:2012pb ; Harlander:2016hcx .

### v2

@@ 3 Results
In this section we present our results for the finite-quark-mass effects in the gluon-gluon-fusion channel for various computational setups. Following the Higgs Working Group recommendations LHCHiggsCrossSectionWorkingGroup:2016ypw , our reference computation uses the 5FS, an OS-renormalised top quark with a pole mass of $m_{t}=173.06$ GeV [^2] The Higgs-Working-Group recommends using an OS-mass of $m_{t}=172.4$ GeV. The different value was chosen because of the rationalisation procedure described in sec. 2. We have verified, that choosing a different mass has a negligible impact on the final result. and an $\overline{\text{MS}}$-renormalised bottom mass of $\overline{m}_{b}(\overline{m}_{b})=4.18$ GeV. We then explore the impact of alternative mass-renormalisation schemes for both the top and the bottom quark. In addition, we perform the computation in the 4FS to examine if the flavour schemes are compatible.
@@
Table 3: Effect of the finite top-quark mass on the gluon-gluon fusion cross section for two different computational setups. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set in the 5FS and the NNPDF31_nnlo_as_0118_nf_4 PDF set in the 4FS. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation.
@@
Tab. 5 presents the results for the top-bottom interference contribution to the cross section $\sigma_{t\times b}$. As already noticed in ref. Czakon:2023kqm , the interference converges quite badly if the bottom-quark mass is OS-renormalised, as indicated by the alternating perturbative corrections of a similar order of magnitude (see third column of tab. 5). Furthermore, scale uncertainties fail to correctly estimate the effects of higher orders and actually increase from NLO to NNLO, reaching 15%. However, results for $\overline{\text{MS}}$-renormalised bottom-quark masses (second column) show a much better perturbative convergence and additionally, we observe a decrease in scale uncertainties to below 8% [^3] Adapting a mixed renormalisation scheme, in which only the Yukawa coupling is renormalised in the $\overline{\text{MS}}$ scheme, while the mass in the propagators is kept in the OS scheme, also results in a better convergence and reduced scale uncertainties as noted in ref. Czakon:2023kqm . Without the inclusion of electroweak corrections this mixed approach is valid, but if these effects are taken into account as well, the Higgs vacuum expectation value acquires an additional finite renormalisation that propagates to the masses of gauge bosons. Since the masses of the gauge bosons are defined in the on-shell scheme by default, the mixed approach is inconsistent. Ignoring this issue, we note that the main improvements to the perturbation series of our problem are already captured by the mixed scheme.. For the top-quark mass, the renormalisation scheme seems to have very little impact on the final result. Lastly, the NNLO result in the 4FS is 5.2% lower than in the 5FS and agrees well within scale uncertainties. The magnitude of this decrease is of a similar size as observed in HEFT (see tab. 2).
Table 5: Top-bottom interference contribution to the gluon-gluon fusion cross section for various computational setups. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set in the 5FS and the NNPDF31_nnlo_as_0118_nf_4 PDF set in the 4FS. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation.
@@
Figure 6: Interference contribution to the Higgs-$p_{T}$ (left) and Higgs-rapidity (right) distribution. Compared are the 5FS results for an $\overline{\text{MS}}$- and OS-renormalised bottom-quark mass. The lower panel shows the ratio to the distribution including finite bottom- and top-quark-mass effects, i.e. finite top-quark masses and the interference to the bottom quark, at NNLO. Coloured bands indicate the envelope of the seven-point scale variation and the error bars represent the Monte Carlo uncertainties.
@@
Figure 7: Higgs-$p_{T}$ (left) and Higgs-rapidity (right) distribution for the gluon-gluon fusion channel. Compared are 5FS results obtained in HEFT with those including finite-quark-mass effects in the square of the top-quark-loop amplitude and in the interference of the top-quark-loop and bottom-quark-loop amplitudes. The bottom-quark mass is renormalised in the $\overline{\text{MS}}$ scheme, while the top-quark mass is renormalised in the OS scheme.
For the Higgs-rapidity distribution in fig. 7 we also observe that the scale-uncertainty bands of the different perturbative orders do not overlap. For both full QCD and HEFT, the difference between NLO and NNLO amounts to a rapidity independent upwards shift of about 20%. Similarly, finite quark masses also only shift the distribution, this time by around $-4\%$, an effect which is directly derived from the interference contribution in fig. 6.
@@ 4 Conclusions
In this work, we studied the impact of using variable flavour number and different mass-renormalisation schemes on the Higgs production cross section in the gluon-gluon fusion channel. We found that the two different flavour schemes are in excellent agreement for the top-bottom interference contribution, finally justifying our treatment of the bottom-quark mass in the 5FS. The total cross section differs by around 4% between different FSs and is consistent within scale uncertainties. Generally speaking, it is advisable and more common to use the 5FS for LHC phenomenology because it automatically resums logarithms of the form $\log(m_{b}^{2}/s)$. Our computation in two different renormalisation schemes showed that renormalising the bottom-quark mass in the $\overline{\text{MS}}$ scheme results in significantly smaller scale uncertainties and displays a better perturbative convergence. The renormalisation of the top-quark mass on the other hand does not impact the cross section significantly. We therefore recommend using the 5FS, an $\overline{\text{MS}}$-renormalised bottom mass and an OS-renormalised top mass going forward. The cross section for this set-up reads [^5] The $\mathrm{N}^{3}\mathrm{LO}$ HEFT cross section was computed using SusHi Harlander:2012pb ; Harlander:2016hcx .
@@ Appendix A Comparison to Higgs+jet analyses
Appendix B Alternative HEFT prescription
Instead of defining the HEFT framework in terms of the LO $K$-factor (see eq.2), some studies like ref. Anastasiou:2016cez additionally consider an NLO rescaling prescription defined by
$\sigma^{\text{N}^{n}\text{LO}}_{\text{HEFT}}=\frac{\sigma^{\text{NLO}}}{\sigma^{\text{NLO}}_{\text{HTL}}}\sigma_{\text{HTL}}^{\text{N}^{n}\text{LO}}\approx 1.056\times\sigma^{\text{N}^{n}\text{LO}}_{\text{HTL}}.$ (16)
The rescaling factor $\sigma^{\text{NLO}}/\sigma_{\text{HTL}}^{\text{NLO}}$ is evaluated at the central scale only, since a scale dependent rescaling factor would result in a renormalisation-scheme dependent HEFT cross section.
The effect of finite top-quark masses with this HEFT prescription is displayed in tab. 6. We see that the effect of the finite top-quark-masses is of similar magnitude—though slightly larger—than in our original rescaling method. Hence, both prescriptions provide accurate approximations for finite top-quark mass effects.
Table 6: Effect of the finite top-quark mass on the gluon-gluon fusion cross section for two different computational setups. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set in the 5FS and the NNPDF31_nnlo_as_0118_nf_4 PDF set in the 4FS. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation.
Order $(\sigma_{t}-\sigma_{\text{HEFT}})$ [pb]
$\sqrt{s}=13$ TeV
5FS 4FS
$m_{t}=173.06$ GeV $m_{t}=173.06$ GeV
$\overline{m}_{b}(\overline{m}_{b})=4.18$ GeV
LO $0.13^{+0.04}_{-0.03}$ $0.12^{+0.03}_{-0.02}$
NLO $0^{+0.08}_{-0.17}$ $0^{+0.05}_{-0.13}$
NNLO $0.23^{+0.16}_{-0.07}$ $0.18^{+0.13}_{-0.06}$